OpenStudy (caitlinnr14):

HELP WILL GIVE MEDAL FOR BEST ANSWER ]Identify the correct slope and y intercept of the equation x - 4y = 4. slope = 4; y intercept at (0, -1) slope = one fourth; y intercept at (0, 1) slope = one fourth; y intercept at (0, -1) slope = -4; y intercept at (0, 1)

OpenStudy (anonymous):

rearrange in the form of y = mx+c y=x/4-1 m = 1/4 c=-1 so option 3rd